OpenStudy (jalil.h):

What is the product of -4 1/2 and 2 1/2 Enter your answer in the box as a mixed number in simplest form.

OpenStudy (jalil.h):

@AaronAndyson @Rosie62122020

OpenStudy (harman.singh):

(-4 1/2) * (2 1/2) = (-9/2) * (5/2) = (-4.5) * (2.5) = -11.25 or -45/4 In mixed number form, it equals to 11 1/4

OpenStudy (harman.singh):

Source: www.webmath.com

OpenStudy (harman.singh):

Correction: I made a mistake in the final answer. It should be -11 1/4 . Forgot to add the negative sign before